Find the equation of a parabola with the vertex at (0,0) and the focus at (0,-1).
Answer by DSMLMD(16)

Vertex at (0,0) and focus at (0,-1)


vertex:
(h,k)
(0,0)


focus point:
(0,-1)
(0,p)


we get p = -1


so because the point of focus located in y-coordinate, so the parabola equation is:


(x^2 - h) = 4p(y - k)
(x^2 - 0) = 4(-1)(y - 0)
x^2 = -4y


The parabola equation is x^2 = -4y